Solve.
(7+u)(3u-1)=0
(If there is more than one solution, separate them with commas.)

Answer by Alan3354(69443)

(7+u)(3u-1)=0
-----------
For the product to be zero, one of the factors has to be zero.
7+u = 0
u = -7
------
3u-1 = 0
u = 1/3